OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (anonymous):

It's The second one ..... y-6=-3(x+1) :)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

Oh so welcome :)! my pleasure

OpenStudy (anonymous):

@xXNoTempestXx would you help me out on one more?

OpenStudy (anonymous):

Of Course ! Go ahead

OpenStudy (anonymous):

Given the equation y - 3 = one half(x + 6) in point-slope form, identify the equation of the same line in standard form. x - 2y = -12 y = one halfx y = one halfx + 6 y = one halfx + 9

OpenStudy (anonymous):

The formula for standard form is Ax+By=C. So, the only possible answer is the first one. I did the work though and it works ! :)

OpenStudy (anonymous):

Okay! Your literally a lifesaver thank you(:

OpenStudy (anonymous):

That means alot ! Haha, no problem !